did i do this right?
rounded 2 decimal places the approximate x intercept of the equation x^3+14.2x^2-4.8x-12.4=0
in standard window
i got 53.44??????????????

Answer by venugopalramana(3286)

NO ANSWER IS ABOUT X=1.07184…CHECK….
X^3= 1.231
14.2*X^2= 16.314
-4.8X= -5.145
-12.4= -12.400
TOTAL= 0.000